OpenStudy (anonymous):

find the slope between x=0 and x=3, for the function: -sq(2x^2 -x +4)

OpenStudy (anonymous):

\[-\sqrt{2x ^{2} - x +4}\]

OpenStudy (anonymous):

@satellite73 has this question meaning to u?

OpenStudy (anonymous):

Is the wording confusing?

OpenStudy (anonymous):

Find the average rate of change of the function f over the given interval x = 0 to x = 3 for the above function

OpenStudy (anonymous):

yes it is i dont know what the problem ask i think it must be find the slope at the point x=3

OpenStudy (anonymous):

ahhhh!!!!

OpenStudy (anonymous):

it means \[\frac{f(3)-f(0)}{3-0}\]

OpenStudy (anonymous):

you got it from there?

OpenStudy (anonymous):

right, could you please do it out for me? the answer they're giving is -3.4801

OpenStudy (anonymous):

and I'm getting -.07863, and I don't know why

OpenStudy (anonymous):

\[f(0)=-\sqrt{4}=-2\] right?

OpenStudy (anonymous):

and \(f(3)=-\sqrt{19}\) i think

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so my answer would be \[\frac{-\sqrt{19}+2}{3}\]

OpenStudy (anonymous):

i have no idea what this looks like as a decimal, but i could check

OpenStudy (anonymous):

ok, that's the same as mine! thank you, I was worried I was doing it wrong

OpenStudy (anonymous):

yw